Christine Haining Pharmacy 487 October 27, 1991

Taxol and Cardiotoxicity

Taxol is a novel antineoplastic drug isolated from the Western Yew, taxus brevifolia. It is novel because of its mechanism of action. Unlike other antimicrotubular agents (e.g., vinca alkaloids, colchicine) which induce microtubular disassembly, taxol promotes microtubule assembly and stability. This stability blocks the dynamic process of microtubule reorganization that normally occurs in the intact cell, and thus cellular replication is inhibited (usually in the G2/M phase). Indications for taxol include: melanoma, non-small cell lung, ovarian, breast, head and neck, colon, and pancreatic carcinomas. With so many indications people are anxious to begin widespread trials. However, taxol is, of course, not free of side effects. One of these side effects is cardiotoxicity. This paper will explore the types of cardiac events, who typically has problems, and who should be monitored for these problems.

It is not presently clear if the cardiotoxicity is caused by taxol or the vehicle (cremophor).^{1} Cremophor has been implicated as the cause of cardiotoxicity, and its side effects when given alone include: vasodilation, labored breath, lethargy, and presumed hypotension.^{3} However, when taxol is given (with cremophor) alone or in a combination with cisplatin the following cardiotoxicities were observed: ventricular arrhythmias, bradycardia, several degrees of atrioventricular conduction blocks, bundle branch block, cardiac ischemia, and myocardial infarction.^{1} It is interesting to note that neither ventricular tachycardia (VT) nor bundle branch blocks had been previously reported with taxol alone, but surfaced as a problem when taxol was combined with cisplatin.^{2}

One study found that 29% of the patients who had no prior history of cardiovasculature problems showed EKG changes.^{1} However, most of the cardiac events observed were asymptomatic bradycardias^{1} when taxol was used alone, and asymptomatic VT when taxol and cisplatin were used together.^{1} In a few cases the bradycardia or VT became symptomatic and required stopping the infusion. EKG abnormalities generally resolved quickly and many of the patients went on to get several more doses of taxol without cardiac event.

Patients with a prior history of hypertension, artherosclerosis, and documented cardiac arrhythmias of any sort are at a much higher risk of serious cardiac events.^{1} For example, a man with a history of hypertension and hypercholesterolemia during the 19th hour of a 24 hour infusion developed complete heart block and died.^{1} There are several other cases where patients with cardiac histories developed cardiac events, such as atrioventricular block, that resolved several hours after discontinuation of the drug.^{1}

It would seem prudent, especially with the onset of broad clinical trials in the near future, to have an EKG running with each patients first infusion of taxol. Patients with a cardiac history should be monitored with every infusion. Again, when cisplatin and taxol are given together all patients with cardiac histories should be monitored every time. The decision as to whether or not monitoring should be performed on patients who have no prior history of cardiac problems is a bit murkier. This is because patients who developed VT when given cisplatin and taxol never develop VT on the first dose. VT was noted either after the first infusion or in the latter half of four courses.^{2} Because VT became serious enough to stop infusion for several patients in this study, it would be prudent to monitor everyone until more data can give a conclusive answer as to EKG monitoring with this combination of drugs.

Finally, it should be “noted that the conduction disturbances do not always recur with repeat courses of taxol, so it may be very reasonable to continue therapy despite conduction abnormalities, depending on individual circumstances.”^{1}

REFERENCES

1. Rowinsky, Eric K., McGuire, W. P., Guarnieri, T., et al: Cardiac disturbances during the administration of taxol. J Clin Oncol 9:1704-1712, 1991

2. Rowinsky, Eric K., Gilbert, M. R., McGuire, W. P., et al: Sequences of taxol and cisplatin: a phase I and pharmacologic study. J Clin Oncol 9:1692-1703, 1991

3. Donehower, Ross C., Rowinsky, Eric K., Grochow, L. B., et al: Phase I trial of taxol (NSC 125973) in patients with advanced cancer. Cancer Treat Rep 71:1171-1177, 1987